OpenStudy (anonymous):

Use the explicit formula an = a1 + (n - 1) d to find the 500th term of the sequence below. 24, 30, 36, 42, 48, ..

OpenStudy (mathmath333):

\(\large \tt \begin{align} \color{black}{a_1=24\\ d=30-24=36-30=6\\ }\end{align}\)

Nnesha (nnesha):

an mean next term